# Augmented Lagrangian Method based adjoint space framework for sparse reconstruction of acoustic source with boundary measurements

### v1

@@ Abstract
We propose a semismooth Newton-based augmented Lagrangian method for reconstructing sparse sources in inverse acoustic scattering problems. The semismooth Newton method can be iterated in the space of measurements instead of the unknown source to be reconstructed. It is highly efficient, especially when the measurement data is much less than the acoustic source. The source can be calculated from Fenchel-Rockafellar duality theory. We can obtain lots of acceleration and leverage the computational cost. The numerical examples show the high efficiency of the proposed semismooth Newton-based methods.
@@ 1 Introduction
Inverse acoustic scattering is crucial in various applications, including sonar imaging, oil prospecting, and non-destructive testing, among others [17]. In this paper, we focus on the reconstruction of acoustic sources, which is a typical inverse acoustic scattering problem. It is well-known that the inverse source problem lacks uniqueness [8, 30] due to the presence of non-radiating sources, resulting in high ill-posedness. Appropriate regularization that incorporates prior information can leverage the ill-posedness, such as $L^{2}$ regularization [19]. For convenience, we assume that $\Omega\subseteq\mathbb{R}^{d}$ with $d=2$ or $d=3$ is a bounded and compact domain with boundary $\partial\Omega$ of class $C^{3}$ and contains the sources. Sparse regularization is a popular choice for signal processing, including $L^{1}$ regularization and total variation, which can promote sparsity for the solutions or gradients. However, for acoustic scattering involving the Helmholtz equation, which is an infinite-dimensional problem, the existence of a reconstructed solution in $L^{1}(\Omega)$ space can not be guaranteed, due to the lack of weak completeness in $L^{1}(\Omega)$ (see Chapter 4 of [10]). In [9, 15, 16], a larger space, the Radon measure space $\mathcal{M}(\Omega)$, is introduced for inverse problems and sparse optimal control. We also refer to [2] for inverse problems, including electrical impedance tomography with finite measurements and the linearized and locally optimized strategy and algorithms for sparse point acoustic sources [1].
It is known that $\mathcal{M}(\Omega)$ is a Banach space and can be characterized by its dual space $C_{0}(\Omega)$ through the Riesz representation theorem (see Chapter 4 of [10]),
@@
This is also equivalent to $\mathcal{M}(\Omega)=C_{0}(\Omega)^{\prime}$, which means that $\mathcal{M}(\Omega)$ is weakly compact by the Banach-Alaoglu theorem since $C_{0}(\Omega)$ is a separable Banach space [10]. The existence of the reconstructed sparse solution in the Radon measure space $\mathcal{M}(\Omega)$ can be guaranteed. Moreover, since $L^{1}(\Omega)$ can be embedded in $\mathcal{M}(\Omega)$, the regularization with $\mathcal{M}(\Omega)$ can also promote sparsity. Inspired by the developments of the Radon measure space regularization [15, 16], [32] considered the sparse reconstruction of an acoustic source in the Radon measure space $\mathcal{M}(\Omega)$. Full scattering data in $\Omega$ is employed, and semismooth Newton is developed for reconstruction algorithms [32].
@@
There are several iterative methods for inverse scattering with smooth functionals such as $L^{2}$ regularization, including the gradient descent method [47], recursive linearization [5, 6], Newton-type methods, or distorted Born iteration [39, 11, 24, 29, 26, 27]. For nonsmooth regularization, the primal-dual method for the total variation regularized inverse medium scattering problem is developed in [12, 13, 14]. Let us turn to a discrete setting for discussing the efficiency. Suppose that $M$ is the number of unknowns modeling the scatter and $N$ is the number of measurements. It is claimed that “For most practical situations, the number of unknowns is much larger than the number of measurements, i.e., $N\gg M$” in [11]. It was further pointed out that $M\propto\sqrt{N}$ generally for the two-dimensional case [29]. Currently, these Newton-type or distorted Born-type methods are iterated directly in the domain of the unknowns, which means that a linear system such as $Ax=b$ with $A\in\mathbb{C}^{N\times N}$ has to be solved during each Newton update. It can be much more efficient if each Newton update can be carried out in the measurement domain by solving a linear system such as $A_{b}z=c$ with $A_{b}\in\mathbb{C}^{M\times M}$. One then recovers the $N$ unknowns through the $M$ computed dual variables by the optimality conditions directly. This technique can reduce a lot of computational cost.
@@
In this paper, we build a framework that the Newton or semismooth Newton executes on the measurement domain. We call the corresponding solution with the same dimensions of the measurements a “dual solution” or “adjoint field” [21]. We then recover the unknown acoustic source through Fenchel–Rockafellar duality theory. Our contributions can be summarized as follows. First, we develop a semismooth Newton-based augmented Lagrangian method, where we can employ semismooth Newton methods in the adjoint space and the linear update of the Newton method to solve problems with the same dimensions as the boundary measurements. Such ideas for inverse electromagnetic scattering can be found in [21, 22], where incremental iterative methods are developed that involve a linear equation with the same dimensions as the measurements. Compared to [21, 22], our framework can handle nonsmooth regularization, whereas $L^{2}$ Tikhonov regularization is employed in [21, 22]. The semismooth Newton method thus can be highly efficient, while the boundary measurements can be much less than the unknown acoustic source to be reconstructed. We would like to point out that a similar dual framework is studied for statistical or signal data processing [37, 48]. Second, we also design a semismooth Newton method with the Moreau–Yosida regularization [16, section 3.1], [28, Chapter 9] or [32] for directly reconstructing the acoustic source in case the scale of the measurement data is comparable with the unknowns.
@@ 2 Inverse source problem with boundary measurements
Reconstructing a sparse source $f$ in the Radon measure space $\mathcal{M}(\Omega)$ for a given scattered field on the boundary $\partial\Omega$.
The direct scattering problem in the frequency domain with an inhomogeneous medium in $\mathbb{R}^{d}$ with $d=2$ or $d=3$ is as follows
@@
where $\mu\in\mathcal{M}(\Omega)$ is a Radon measure and $n(x)$ is the refraction index. We refer to Figure 1 for an illustration. Henceforth, we assume $n(x)$ is real, smooth, and compactly supported, i.e., $\Im n=0$ and there exists a bounded domain $\Omega\subseteq\mathbb{R}^{d}$ such that $\text{supp}(n(x)-1)\Subset\Omega$. Throughout this paper, we assume $\mu$ is a real Radon measure, which is reasonable in physics and is also compactly supported. We assume the bounded domain $\Omega$ is large enough such that $\mu$ is also compactly supported in $\Omega$, i.e., $\text{supp}(\mu)\Subset\Omega$. The Helmholtz equation (2.1) can also be recovered with $n(x)\equiv 1$.
The solution of (2.1) is defined in “very weak” sense [32, Definition 2.1] with test functions in $C^{2,\alpha}$. Here, for convenience, we directly work with the following volume potential representations,
@@
where the inhomogeneous and background Green’s function $G(x,y)$ is defined as the radiating solution [18]
@@
Denoting m(x)=1n(x)m(x)=1-n(x), we thus can construct $G(x,y)$ by the Lippmann-Schwinger integral equation [18]
@@
Figure 1: Illustration of the acoustic source scattering: the red squares represent the receivers for recording the scattering wave, and the trapezoid AA represents the acoustic sources.
@@
$\min_{\mu}\frac{1}{2}\|\mathcal{V}\mu-u_{b}\|_{L^{s}(\partial\Omega)}^{2}+\alpha\|\mu\|_{\mathcal{M}(\Omega)},\quad\forall s\begin{cases}\text{such that }\ p\leq s<\infty,\ \ &d=2,\\ \text{such that }p\leq s<2,\ \ &d=3,\end{cases}$ (2.8)
@@
The proof is similar to [16, Proposition 2.2] and [9]. Since the energy in (2.8) is $\frac{1}{2}\|u_{b}\|_{L^{s}(\partial\Omega)}^{2}$ while $\mu=0$. Thus we can find a minimizing sequence $\{\mu_{n}\}$ in $\mathcal{M}(\Omega)$ which are bounded by $\frac{1}{2\alpha}\|u_{b}\|_{L^{s}(\partial\Omega)}^{2}$. Since $\mathcal{M}(\Omega)$ is a weakly sequentially compact [10] (see Chapter 4), there exists a weakly convergent subsequence $\mu_{n,k}$ converging to $\mu^{*}\in\mathcal{M}(\Omega)$ weakly.
Denoting $u_{n,k}=\mathcal{V}(\mu_{n,k})$, we see $u_{n,k}\in W^{1,p}(\Omega)$ with $p<\frac{d}{d-1}$. By Proposition 1, $u_{n,k}$ weakly converges to $\mathcal{V}(\mu^{*})$ and $\mathcal{V}(\mu^{*})$ is a solution (2.1) within definition of the very weak solution [32, Definition 2.1] defined as the dual of $C^{2,\alpha}$. By the weak lower semicontinuity of the norms in $L^{2}(\Omega)$ and $\mathcal{M}(\Omega)$, we conclude that $\mu^{*}$ is a minimizer of (2.8) and the existence follows. ∎
For the non-smooth minimization problem (2.8), it is convenient to consider the predual problem under the powerful Fenchel–Rockafellar duality theory; see [9, 28, 15, 16] for its various applications in inverse problems and optimal control problems. The semismooth Newton method can be employed to compute dual problems efficiently. However, the problem (2.8) is with a complex-valued function. For the application of Fenchel–Rockafellar duality theory, we need to reformulate the complex-valued operators and functions in terms of real matrix operators and real vector functions. Let us denote
@@
where $\mathcal{V}_{R}=\Re(\mathcal{V})$, $\mathcal{V}_{I}=\Im(\mathcal{V})$. We still employ $\mu$ and $u_{b}$ as their complex realization.
@@ 3 Semismooth Newton-based Augmented Lagrangian Method
Henceforth, we consider the discrete setting of the functional in (2.8) and denote $\mathcal{V}_{b}=V|_{\partial\Omega}$, i.e, the trace of the volume potential after complex realization. For the discretization of $\mathcal{V}_{b}$, we employ the framework in [31], which is implemented in [12, 13] with publicly available code. The inhomogeneous medium can be solved by the Lippmann-Schwinger equation; we will leave some comments in section 5. Since $\mathcal{V}_{b}$ may lack of injectivity, for the convenience of the semismooth Newton methods along with optimization framework, we introduce an extra $L^{2}$ regularization, and work in $L^{2}(\partial\Omega)$ instead of $L^{s}(\partial\Omega)$ in discrete settings,
@@
Let us first present the general Fenchel-Rockafellar duality theory [28, Theorem 4.34] or [7, Theorem 15.23]. Let $f$, $g$ be convex, proper, lower-semicontinuous functions in Hilbert spaces $Y$ and $X$, and $\Lambda:X\rightarrow Y$ is a linear, bounded, and continuous operator. Then, under certain conditions (such as $0\in\text{sri}(\operatorname{dom}f-\Lambda(\operatorname{dom}g))$ [7, Theorem 15.23] with “sri” denoting strong relative interior and $\operatorname{dom}$ denoting the definition domain), the dual problem of the following primal minimization problem
@@
Here $x^{*}$ and $y^{*}$ are the solutions of the primal problem (3.3) and dual problem (3.4) respectively. $\Lambda^{*}:X\rightarrow Y$ is the adjoint operator of $\Lambda$. The optimality conditions (3.5) can shuttle solutions between the primal and the dual solutions.
@@ 3.1 Semismooth Newton-based augmented Lagrangian framework in the dual space
and $\text{sign}(\cdot)$ is the sign function. It will turn out that the explicit formulation of $p^{*}$ is not needed later. By Fenchel–Rockafellar duality, we obtain the dual problem of (3.6)
@@
which is equivalent to $\max_{y\in L^{2}(\partial\Omega)}(-p^{*}(-\mathcal{V}_{b}^{T}y)-h^{*}(y))$ by (3.4). Furthermore, letting $\mathcal{V}_{b}^{T}y=z$, the above minimization problem becomes the following linear constrained optimization problem
@@
The Augmented Lagrangian method (shortened ALM) differs from ADMM (alternating direction method of multipliers), as ALM solves $(y,z)$ variables together, whereas ADMM solves $y$ and $z$ separately. ALM can be obtained from the proximal point algorithm [41] to the dual problem [42], and increasing the step sizes $\sigma_{k}$ can obtain superlinear convergence for certain functions. For the developments of ALM, we refer to [35, 38] and [33, 34].
Semismooth Newton methods [28, 49] are widely employed for the nonlinear and nonsmooth updates of ALM [37, 50] and [44, 25, 36]. For its application in total variation (or total generalized variation) regularized imaging processing problems, we refer to [45, 46]. Let us turn to semismooth Newton methods for solving the coupled nonlinear system of $(y,z)$ in (3.14). The first order optimality conditions on $y^{k+1}$ and $z^{k+1}$ respectively of (3.14) tell that
@@
Usually, for discretizations, the dimensions of $L^{2}(\partial\Omega)$ is much less than the dimensions of $L^{2}(\Omega)$ [11]. It can be much more efficient if only the Newton update of $y^{k+1}$ is solved. Letting us solving $z^{k+1}$ with $y^{k+1}$ first. By $(\sigma_{k}I+\partial p^{*})(z^{k+1})\ni-\sigma\mathcal{V}_{b}^{T}y^{k+1}-\lambda^{k}$, we now solve $z^{k+1}$ from (3.15)
@@
Substituting the above formula which is $\sigma_{k}z^{k+1}$ actually into (3.15), we arrive at
@@
This leads to the following nonlinear equation of $y^{k+1}$ on $L^{2}(\partial\Omega)$ finally
@@
Note that $y\in L^{2}(\partial\Omega)$ and the nonlinear equation (3.20) is on the boundary $\partial\Omega$, which can be of much smaller scale compared to the source $\mu$ to be reconstructed. Since $(I+\sigma\partial p)^{-1}(\cdot)$ is a semismooth function for any positive $\sigma$, we have the following lemma for the semismooth Newton derivative of $\mathcal{F}$.
@@
Now, introducing $l(y):=(I+\sigma_{k}\partial p)^{-1}(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)$, we can write its expression component-wisely
@@
It can be checked that $l(y)$ is a $PC^{1}$ function (piecewise differentiable function) [43, Proposition 4.3.1]. Since $l(y)$ is continuous differentiable on $y_{i}$ when $|(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)_{i}|>\sigma_{k}\alpha$, we have $(\partial l)(z)]_{i}=(-\frac{\sigma}{1+\sigma_{k}\alpha_{0}}\mathcal{V}_{b}(z))_{i}$. While $|(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)_{i}|<\sigma_{k}\alpha$, we obtain $(\partial l)(z)]_{i}=0$. For the case $|(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)_{i}|=\sigma_{k}\alpha$, by [43, Theorem 4.3.1], we have $s(\frac{-\sigma_{k}}{1+\sigma_{k}\alpha_{0}}\mathcal{V}_{b}(z))_{i}\in[(\partial l)(z)]_{i}$ for any $s\in[0,1]$. Since the soft thresholding operator $\mathcal{S}_{\frac{\sigma_{k}\alpha}{1+\sigma_{k}\alpha_{0}}}$ is anisotropic on $y$. We finally arrive at the semismooth Newton derivative $\partial F$ as in (3.21). ∎
The Newton update for solving $y^{k+1}$ becomes that for $l=0,2,\ldots,$ with $y^{0}=y^{k}$
@@
It can be seen that from (3.21), the semismooth Newton derivative is a positive definite and bounded operator. Thus for the Newton update in (3.26) and any $\mathcal{N}(y^{l})\in\partial F(y^{l})$, we have $\mathcal{N}(y^{l})\succ I$ and $\mathcal{N}(y^{l})^{-1}$ is uniformly bounded. The Newton update (3.26) is well-defined.
@@
Since Newton methods, including semismooth Newton methods, are always locally convergent. Line search is usually employed for globalization and convergence. We use the following Armijo line search strategy. By defining $d^{l}=y^{l+1}-y^{l}$, we proceed with an aggressive Armijo-type line search involving parameters $\beta^{t}>0$, $t=0,1,2,\ldots$, $\beta\in(0,1)$, and $c>0$. The objective is to find the smallest integer $t$ satisfying:

### v3

# An Augmented Lagrangian Method-Based Framework in the Adjoint Space for Sparse Reconstruction of Acoustic Sources

@@ Abstract
We propose a semismooth Newton-based augmented Lagrangian framework for reconstructing sparse sources in inverse acoustic scattering problems. Rather than working in the unknown-source space, our semismooth Newton updates operate in the measurement (adjoint) space, which is especially efficient when the number of measurements is much smaller than the discretized source dimension. The source is then recovered via Fenchel–Rockafellar duality. Our approach substantially accelerates computation and reduces costs. Numerical experiments in two and three dimensions demonstrate the high efficiency of the proposed method.
Key words.
inverse acoustic scattering, inverse source problem, adjoint space method, semismooth Newton method, augmented Lagrangian method, global convergence
@@ 1 Introduction
Inverse acoustic scattering plays a key role in applications such as sonar imaging, oil prospecting, and nondestructive testing [17]. Here we focus on the inverse source problem: reconstructing acoustic sources from boundary measurements. It is well-known that the inverse source problem lacks uniqueness [8, 30] due to the presence of non-radiating sources, resulting in high ill-posedness. Appropriate regularization that incorporates prior information can mitigate the ill-posedness (e.g., $L^{2}$ regularization [19]. For convenience, we assume that $\Omega\subseteq\mathbb{R}^{d}$ with $d=2$ or $d=3$ is a bounded and compact domain with boundary $\partial\Omega$ of class $C^{3}$ and contains the sources. Sparse regularization is a popular choice in signal processing, including $L^{1}$ regularization and total variation, which can promote sparsity in the solutions or their gradients. However, for acoustic scattering involving the Helmholtz equation, which is an infinite-dimensional problem, the existence of a reconstructed solution in the $L^{1}(\Omega)$ space cannot be guaranteed due to the lack of weak completeness in $L^{1}(\Omega)$ (see [10, Chapter 4]). In [9, 15, 16], a larger space, the Radon measure space $\mathcal{M}(\Omega)$, is introduced for inverse problems and sparse optimal control. We also refer to [2] for inverse problems, including electrical impedance tomography with finite measurements and a linearized and locally optimized strategy and algorithms for sparse point acoustic sources [1].
It is known that $\mathcal{M}(\Omega)$ is a Banach space and can be characterized by its dual space $C_{0}(\Omega)$ through the Riesz representation theorem (see [10, Chapter 4]),
@@
This is also equivalent to $\mathcal{M}(\Omega)=C_{0}(\Omega)^{\prime}$, which means that $\mathcal{M}(\Omega)$ is weakly compact by the Banach-Alaoglu theorem since $C_{0}(\Omega)$ is a separable Banach space [10]. The existence of a reconstructed sparse solution in the Radon measure space $\mathcal{M}(\Omega)$ can be guaranteed. Moreover, since $L^{1}(\Omega)$ can be embedded in $\mathcal{M}(\Omega)$, regularization in $\mathcal{M}(\Omega)$ can also promote sparsity. Inspired by developments in the Radon measure space regularization [15, 16], [32] considered the sparse reconstruction of an acoustic source in the Radon measure space $\mathcal{M}(\Omega)$. The complete scattering data in $\Omega$ is employed, and a semismooth Newton method is developed for the reconstruction algorithms [32].
@@
There are several iterative methods for inverse scattering with smooth functionals such as $L^{2}$ regularization, including the gradient descent method [47], recursive linearization [5, 6], Newton-type methods, and distorted Born iteration [39, 11, 24, 29, 26, 27]. For nonsmooth regularization, a primal-dual method for the total variation regularized inverse medium scattering problem is developed in [12, 13, 14]. Let us turn to a discrete setting to discuss the efficiency. Suppose that $N$ is the number of unknowns modeling the scatterer and $M$ is the number of measurements. As noted in [11], “For most practical situations, the number of unknowns is much larger than the number of measurements, i.e., $N\gg M$”. In [29], it was also pointed out that $M\propto\sqrt{N}$ generally holds for the two-dimensional case. Currently, these Newton-type or distorted Born-type methods are iterated directly in the domain of unknowns, which means that a linear system such as $Ax=b$ with $A\in\mathbb{C}^{N\times N}$ must be solved during each Newton update. It can be much more efficient if each Newton update can be performed in the measurement domain by solving a linear system such as $A_{b}z=c$ with $A_{b}\in\mathbb{C}^{M\times M}$. The $N$ unknowns are then recovered from the $M$ computed dual variables directly via the optimality conditions. This technique can significantly reduce the computational cost.
@@
In this paper, we build a framework in which the Newton or semismooth Newton operates in the measurement domain. We call the corresponding solution (with the same dimension as the measurements) a “dual solution” or “adjoint field” [21]. We then recover the unknown acoustic source through Fenchel–Rockafellar duality theory. Our contributions can be summarized as follows. First, we develop a semismooth Newton-based augmented Lagrangian method in which semismooth Newton updates are performed in the adjoint space. It enables us to solve linear systems with the same dimension as the boundary measurements. Such ideas for inverse electromagnetic scattering can be found in [21, 22], where incremental iterative methods are developed that involve a linear equation with the same dimensions as the measurements. Compared to [21, 22], our framework can handle nonsmooth regularization, whereas $L^{2}$ Tikhonov regularization is employed in [21, 22]. The semismooth Newton method can thus be highly efficient when the number of boundary measurements is much smaller than the number of unknowns in the acoustic source to be reconstructed. We would like to point out that a similar dual framework is studied in statistical or signal data processing [37, 48]. Second, we also design a semismooth Newton method with Moreau-Yosida regularization (see [16, section 3.1], [28, Chapter 9], or [32]) to directly reconstruct the acoustic source in cases where the scale of the measurement data is comparable to that of the unknowns.
@@ 2 Inverse source problem with boundary measurements
Reconstructing a sparse source $\mu$ in the Radon measure space $\mathcal{M}(\Omega)$ from a given scattered field on the boundary $\partial\Omega$.
The direct scattering problem in the frequency domain with an inhomogeneous medium in $\mathbb{R}^{d}$ with $d=2$ or $d=3$ is as follows:
@@
where $\mu\in\mathcal{M}(\Omega)$ is a Radon measure and $n(x)$ is the refractive index. We refer to Figure 1 for an illustration. Henceforth, we assume that $n(x)$ is real, smooth, and compactly supported, i.e., $\Im n=0$ and there exists a bounded domain $\Omega\subseteq\mathbb{R}^{d}$ such that $\text{supp}(n(x)-1)\Subset\Omega$. Throughout this paper, we assume that $\mu$ is a real, compactly supported Radon measure, which is reasonable in physics. We assume that the bounded domain $\Omega$ is large enough such that $\mu$ is also compactly supported in $\Omega$, i.e., $\text{supp}(\mu)\Subset\Omega$. The Helmholtz equation (2.1) can also be considered for the case $n(x)\equiv 1$.
The solution of (2.1) is defined in a “very weak” sense [32, Definition 2.1] using test functions in $C^{2,\alpha}$. Here, for convenience, we directly work with the following volume potential representations,
@@
where the inhomogeneous background Green’s function $G(x,y)$ is defined as the radiating solution [18]
@@
Denoting m(x)=1n(x)m(x)=1-n(x), we can thus construct $G(x,y)$ by the Lippmann-Schwinger integral equation [18]
@@
Figure 1: Illustration of the acoustic source scattering: the red squares represent the receivers for recording the scattered wave, and the trapezoid labeled AA represents the acoustic sources.
@@
$\min_{\mu}J(\mu):=\frac{1}{2}\|\mathcal{V}\mu-u_{b}\|_{L^{s}(\partial\Omega)}^{2}+\alpha\|\mu\|_{\mathcal{M}(\Omega)},\quad\forall s\begin{cases}\text{such that }\ p\leq s<\infty,\ \ &d=2,\\ \text{such that }p\leq s<2,\ \ &d=3,\end{cases}$ (2.8)
@@
The proof is similar to [16, Proposition 2.2] and [9]. Since the energy in (2.8) is $\frac{1}{2}\|u_{b}\|_{L^{s}(\partial\Omega)}^{2}$ when $\mu=0$, so we can find a minimizing sequence $\{\mu_{n}\}_{n}$ in $\mathcal{M}(\Omega)$ that is bounded by $\frac{1}{2\alpha}\|u_{b}\|_{L^{s}(\partial\Omega)}^{2}$. Since $\mathcal{M}(\Omega)$ is a weakly sequentially compact [10] (see Chapter 4), there exists a weakly convergent subsequence $\mu_{n,k}$ converging weakly to $\mu^{*}\in\mathcal{M}(\Omega)$.
Denoting $u_{n,k}=\mathcal{V}(\mu_{n,k})$, we see $u_{n,k}\in W^{1,p}(\Omega)$ with $1\leq p<\frac{d}{d-1}$ with (2.5). By Proposition 1, $u_{n,k}$ weakly converges to $\mathcal{V}(\mu^{*})$, and $\mathcal{V}(\mu^{*})$ is a solution (2.1) in the sense of the very weak solution [32, Definition 2.1], defined as the dual of $C^{2,\alpha}$. By the trace theorem [4, Theorem 5.36] and (2.6), the trace operator is continuous from $W^{1,p}(\Omega)$ to $L^{s}(\partial\Omega)$, so $u_{n,k}|_{\partial\Omega}\rightharpoonup V(\mu^{*})|_{\partial\Omega}$ weakly in $L^{s}(\partial\Omega)$. The data term $\frac{1}{2}\|\cdot-u_{b}\|^{2}_{L^{s}(\partial\Omega)}$ is convex and continuous, hence weakly lower semicontinuous. The regularization term $\alpha\|\cdot\|_{M(\Omega)}$ is a norm, hence weak lower semicontinuous (see [10, Chapter 4]). Therefore,
$J(\mu^{*})\leq\liminf_{n\to\infty}J(\mu_{n})=\inf J,$
so $\mu^{*}$ is a minimizer of (2.8) and the existence follows. ∎
For the nonsmooth minimization problem (2.8), it is convenient to consider the predual problem under the powerful Fenchel–Rockafellar duality theory; see [9, 28, 15, 16] for its various applications in inverse problems and optimal control problems. The semismooth Newton method can be employed to compute dual problems efficiently. However, the problem (2.8) involves a complex-valued function. For the application of Fenchel–Rockafellar duality theory, we need to reformulate complex-valued operators and functions in terms of real matrix operators and real vector functions. Let us denote
@@
where $\mathcal{V}_{R}=\Re(\mathcal{V})$, $\mathcal{V}_{I}=\Im(\mathcal{V})$. We still employ $\mu$ and $u_{b}$ as their complex realizations.
@@ 3 Semismooth Newton-based augmented Lagrangian method
Now, we consider the discrete setting of the functional in (2.8) and denote $\mathcal{V}_{b}=V|_{\partial\Omega}$, i.e., the trace of the volume potential after complex realization. For the discretization of $\mathcal{V}_{b}$, we employ the framework in [31], which is implemented in [12, 13] with publicly available code. The Lippmann-Schwinger equation can be used to solve for the inhomogeneous medium; we provide some comments in Section 5. Since $\mathcal{V}_{b}$ may lack injectivity, for the convenience of the semismooth Newton methods and the associated optimization framework, we introduce an extra $L^{2}$ regularization, and work in $L^{2}(\partial\Omega)$ instead of $L^{s}(\partial\Omega)$ in discrete settings,
@@
Let us first present the general Fenchel-Rockafellar duality theory [28, Theorem 4.34] or [7, Theorem 15.23]. Let $f$, $g$ be convex, proper, lower-semicontinuous functions in Hilbert spaces $Y$ and $X$, and $\Lambda:X\rightarrow Y$ is a linear, bounded, and continuous operator. Then, under certain conditions (such as $0\in\text{sri}(\operatorname{dom}f-\Lambda(\operatorname{dom}g))$ [7, Theorem 15.23], where “sri” denotes strong relative interior and $\operatorname{dom}$ denotes the definition domain), the dual problem of the following primal minimization problem
@@
Here $x^{*}$ and $y^{*}$ are the solutions of the primal problem (3.3) and dual problem (3.4) respectively. $\Lambda^{*}:X\rightarrow Y$ is the adjoint operator of $\Lambda$. The optimality conditions (3.5) can be used to recover solutions between the primal and the dual solutions.
@@ 3.1 Semismooth Newton-based augmented Lagrangian framework in the dual space
and $\text{sign}(\cdot)$ is the sign function. The explicit formulation of $p^{*}$ will not be needed later. By Fenchel–Rockafellar duality, we obtain the dual problem of (3.6)
@@
which is equivalent to $\max_{y\in L^{2}(\partial\Omega)}(-p^{*}(-\mathcal{V}_{b}^{T}y)-h^{*}(y))$ by (3.4). Furthermore, letting $\mathcal{V}_{b}^{T}y=z$, the above minimization problem becomes the following linearly constrained optimization problem
@@
The augmented Lagrangian method (abbreviated as ALM) differs from ADMM (the alternating direction method of multipliers), as ALM solves the $(y,z)$ variables together, whereas ADMM solves $y$ and $z$ separately. ALM can be derived from the proximal point algorithm [41] to the dual problem [42], and increasing the step sizes $\sigma_{k}$ can lead to superlinear convergence for certain functions. For developments of ALM, we refer to [35, 38] and [33, 34].
Semismooth Newton methods [28, 49] are widely employed for the nonlinear and nonsmooth updates of ALM [37, 50] and [44, 25, 36]. For its application in total variation (or total generalized variation) regularized image processing problems, we refer to [45, 46]. We now turn to semismooth Newton methods for solving the coupled nonlinear system of $(y,z)$ in (3.14). The first-order optimality conditions on $y^{k+1}$ and $z^{k+1}$, respectively, in (3.14) are
@@
Usually, for discretizations, the dimensions of $L^{2}(\partial\Omega)$ are much smaller than the dimensions of $L^{2}(\Omega)$ [11]. It can be much more efficient to solve only for the Newton update of $y^{k+1}$. Let us first solve for $z^{k+1}$ in terms of $y^{k+1}$. By $(\sigma_{k}I+\partial p^{*})(z^{k+1})\ni-\sigma\mathcal{V}_{b}^{T}y^{k+1}-\lambda^{k}$, we now solve $z^{k+1}$ from (3.16)
@@
Substituting the above formula (which is actually $\sigma_{k}z^{k+1}$) into (3.15), we arrive at
@@
This finally leads to the following nonlinear equation for $y^{k+1}$ in $L^{2}(\partial\Omega)$
@@
Note that $y\in L^{2}(\partial\Omega)$ and the nonlinear equation (3.20) is on the boundary $\partial\Omega$, which can be of much smaller dimension compared to the source $\mu$ to be reconstructed. Since $(I+\sigma\partial p)^{-1}(\cdot)$ is a semismooth function for any positive $\sigma$, we have the following lemma for the semismooth Newton derivative of $\mathcal{F}$.
@@
Now, introducing $l(y):=(I+\sigma_{k}\partial p)^{-1}(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)$, we can write its expression componentwise
@@
It can be checked that $l(y)$ is a $PC^{1}$ function (piecewise differentiable function) [43, Proposition 4.3.1]. Since $l(y)$ is continuously differentiable on $y_{i}$ when $|(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)_{i}|>\sigma_{k}\alpha$, we have $(\partial l)(z)]_{i}=(-\frac{\sigma}{1+\sigma_{k}\alpha_{0}}\mathcal{V}_{b}(z))_{i}$. Whereas $|(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)_{i}|<\sigma_{k}\alpha$, we obtain $(\partial l)(z)]_{i}=0$. For the case $|(-\lambda-\sigma_{k}\mathcal{V}_{b}^{T}y)_{i}|=\sigma_{k}\alpha$, by [43, Theorem 4.3.1], we have $s(\frac{-\sigma_{k}}{1+\sigma_{k}\alpha_{0}}\mathcal{V}_{b}(z))_{i}\in[(\partial l)(z)]_{i}$ for any $s\in[0,1]$. Since the soft thresholding operator $\mathcal{S}_{\frac{\sigma_{k}\alpha}{1+\sigma_{k}\alpha_{0}}}$ is anisotropic on $y$. We finally arrive at the semismooth Newton derivative $\partial F$ as in (3.21). ∎
The Newton update for solving $y^{k+1}$ are as follows, starting with $y^{0}=y^{k}$ for $l=0,2,\ldots,$
@@
As seen from (3.21), the semismooth Newton derivative is positive definite and bounded. Thus for the Newton update in (3.26) and any $\mathcal{N}(y^{l})\in\partial F(y^{l})$, we have $\mathcal{N}(y^{l})\succ I$ and $\mathcal{N}(y^{l})^{-1}$ is uniformly bounded. The Newton update (3.26) is well-defined.
@@
Since Newton methods, including semismooth Newton methods, are locally convergent, line search is usually employed for globalization and to ensure convergence. We use the following Armijo line search strategy. By defining $d^{l}=y^{l+1}-y^{l}$, we proceed with an aggressive Armijo-type line search involving parameters $\beta^{t}>0$, $t=0,1,2,\ldots$, $\beta\in(0,1)$, and $c>0$. The objective is to find the smallest integer $t$ satisfying: